# Continuum Transformers Perform In-Context Learning by Operator Gradient Descent

### v1

@@ Appendix C In-Context Learning via Gradient Descent Proof
Theorem C.1.
Let $\kappa:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{L}(\mathcal{X})$ be an arbitrary operator-valued kernel and $\mathcal{O}$ be the operator RKHS induced by $\kappa$. Let $\{(f^{(i)},u^{(i)})\}^{n}_{i=1}$ and $L$ be as defined by Equation 11. Let $O_{0}=\mathbf{0}$ and let $O_{\ell}$ denote the operator obtained from the $\ell$-th operator-valued gradient descent sequence of $L$ with respect to $\|\cdot\|_{\mathcal{O}}$ as defined in Equation 25. Then there exist scalar step sizes $r_{0}^{\prime},\ldots,r_{m}^{\prime}$ such that if, for an $m$-layer continuum transformer $\mathcal{T}:=\mathcal{T}_{m}\circ...\circ\mathcal{T}_{0}$, $[\widetilde{H}(U,W)]_{i,j}=\kappa(u^{(i)},w^{(j)})$, $\mathcal{W}_{v,\ell}=\begin{pmatrix}0&0\\ 0&-r^{\prime}_{\ell}I\end{pmatrix}$, $\mathcal{W}_{q,\ell}=I$, and $\mathcal{W}_{k,\ell}=I$ for each $\ell=0,...,m$, then for any $f\in\mathcal{X}$,
@@
From calculations in subsection Section B.1, we know that the $\ell$-th step of gradient descent has the form
@@
for all $j=0,\ldots,n$. In other words, “$u^{(j)}-u^{(j)}_{\ell}$ is equal to the predicted label for $f$, if $f^{(j)}=f$”.
@@
Here, the first line follows from plugging in $\mathcal{W}_{q},\mathcal{W}_{v},\mathcal{W}_{k}$ in Appendix C. The second line follows from Equation 30 and the induction hypothesis. ∎
Appendix D Best Linear Unbiased Predictor
D.1 BLUP Coincides with Bayes Optimal
@@
D.2 In-Context Learning BLUP Proof
Proposition D.2.
Let $F=[f^{(1)},\ldots,f^{(n+1)}],U=[u^{(1)},\ldots,u^{(n+1)}]$. Let $\kappa:\mathcal{X}\times\mathcal{X}\to\mathcal{L}(\mathcal{X})$ be an operator-valued kernel. Assume that $U|F$ is a $\kappa$ Gaussian random variable per Definition 3.2. Let the activation function $\tilde{H}$ of the attention layer be defined as $[\tilde{H}(U,W)]_{i,j}:=\kappa(u^{(i)},w^{(j)})$. Consider the operator gradient descent in Theorem 3.1. Then as the number of layers $m\to\infty$, the continuum transformer’s prediction at layer $m$ approaches the Best Linear Unbiased Predictor that minimizes the in-context loss in Equation 10.
@@ D.2 In-Context Learning BLUP Proof
From the premise of 3.1, we know that $\mathcal{W}_{q,\ell}=I,\mathcal{W}_{k,\ell}=I,\mathcal{W}_{v,\ell}=\begin{bmatrix}0&0\\ 0&-r^{\prime}_{\ell}I\end{bmatrix}$. Set $r_{\ell}=\delta I$ for all $\ell$, where $\delta$ is a positive constant satisfying $\|I-\delta\hat{\mathbb{K}}\|<1$, where the norm is the operator norm. From previous calculations done in Appendix C, we have:
@@
Again, using Appendix C, we have:
@@
Appendix E Optimization Convergence Proof
@@
Theorem E.5.
Suppose E.3 and E.4 hold. Let $f(r,\mathcal{W}_{q},\mathcal{W}_{k}):=L\left(\mathcal{W}_{v,\ell}=\left\{\begin{bmatrix}0&0\\ 0&r_{\ell}\end{bmatrix}\right\}_{\ell=0,\ldots,m},\mathcal{W}_{q,\ell},\mathcal{W}_{k,\ell}\right)$, where $L$ is as defined in Equation 10. Let $\mathcal{S}\subset\mathcal{O}^{m+1}\times\mathcal{O}^{m+1}$ denote the set of $(\mathcal{W}_{q},\mathcal{W}_{k})$ operators with the property that $(\mathcal{W}_{q},\mathcal{W}_{v})\in\mathcal{S}$ if and only if for all $\ell\in\{0,\ldots,m\}$, there exist scalars $b_{\ell},c_{\ell}\in\mathbb{R}$ such that $\mathcal{W}_{q,\ell}=b_{\ell}\Sigma^{-1/2}$ and $\mathcal{W}_{k,\ell}=c_{\ell}\Sigma^{-1/2}$. Then
@@
Clearly, Equation 33 $=-\sum_{\ell=0}^{k}(\partial_{r_{\ell}}L(r(t),\mathcal{W}_{q}(t),\mathcal{W}_{k}(t)))^{2}$. From Proposition E.6, we note that
@@
Proposition E.6.
Suppose $F,U$ satisfy E.3 and $\tilde{H}$ satisfies E.4. Suppose $\mathcal{W}_{q},\mathcal{W}_{k}\in\mathcal{S}$. Fix a layer $j\in\{0,\ldots,m\}$. For any $R\in\mathcal{L}(\mathcal{X})$, let $\mathcal{W}_{q}(R,j)$ denote the collection of operators where $[\mathcal{W}_{q}(R,j)]_{j}:=\mathcal{W}_{q,j}+R$ and $[\mathcal{W}_{q}(R,j)]_{\ell}=\mathcal{W}_{q,\ell}$ for all $\ell\neq j$. Take an arbitrary $R\in\mathcal{L}(\mathcal{X})$. Let
@@
where $R\Sigma^{1/2}$ denotes composition of operators and $Tr$ is the trace of an operator as defined in Definition E.10. Then for any $j\in\{0,\ldots,m\}$,
@@
Suppose $\mathcal{W}_{q}(R)$ be the collection of operators where $[\mathcal{W}_{q}(R)]_{\ell}=\mathcal{W}_{q,\ell}+R$ for all $\ell\in\{0,\ldots,m\}$. Since Proposition E.6 holds for any $j$, and
@@
an immediate consequence of Proposition E.6 is that
@@
Suppose $\bar{Z}_{0}$ is equal to $Z_{0}$ everywhere, except the $(2,n+1)^{th}$ entry, where it is $c$ instead of 00. We compare the dynamics of $Z_{\ell}$ and $\bar{Z}_{\ell}$. Since $F_{0}=\bar{F}_{0}$, we see that $F_{\ell}=\bar{F}_{\ell}$. We claim that $\bar{U}_{\ell}-U_{\ell}=[0,\ldots,0,c]$. We prove this by induction. This is trivially true for $\ell=0$. Suppose this holds for step $\ell$, then at step $\ell+1$:
@@
Replacing $c$ by $u^{(n+1)}$ tells us that $[\bar{Z}_{m+1}]_{2,n+1}=[Z_{m+1}]_{2,n+1}+u^{(n+1)}$. Hence, the loss in Equation 10 can be re-written as
@@
Since $A_{\ell}=0$, an immediate corollary of Appendix E is as follows.
Corollary E.8.
@@
The rest of the proof for Proposition E.6 has the following outline:
@@
Using Lemma E.11, we can write the in-context loss in Appendix E as
@@
Since $(I-M)$ is a self-adjoint operator,
@@
where the last equality follows from E.3 and the linearity and cyclic property of trace.
Let $\mathcal{U}$ be a uniformly randomly sampled unitary operator. Let $\mathcal{U}_{\Sigma}=\Sigma^{1/2}\mathcal{U}\Sigma^{-1/2}$. Using E.3, $\Sigma^{1/2}\mathcal{M}\Sigma^{-1/2}F\stackrel{{\scriptstyle d}}{{=}}F$, we see that
@@
where the last equality uses the assumption $\mathbb{K}(\mathcal{U}_{\Sigma}F_{0})=\mathbb{K}(F_{0})$ from E.3.
@@
where the last equality follows from the rotational invariance of $\tilde{H}$ from E.4.
@@
where the first equality follows from Equation 43 and Equation 44 and the second equality follows from E.4.
@@
where the second equality follows from Equation 43, E.4 and Appendix E.
@@
Continuing from Appendix E, we get
@@
Appendix F Best Linear Unbiased Prediction Additional Experimental Results
@@
Appendix G Optimization Experiment Details
G.1 Optimization Kernel Assumptions
@@
G.2 Optimization Experiment Hyperparameters

### v2

@@ Appendix E In-Context Learning via Gradient Descent Proof
Theorem E.1.
Let $\kappa:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{L}(\mathcal{X})$ be an arbitrary operator-valued kernel and $\mathcal{O}$ be the operator RKHS induced by $\kappa$. Let $\{(f^{(i)},u^{(i)})\}^{n}_{i=1}$ and $L(O):=\sum_{i=1}^{n}\|u^{(i)}-Of^{(i)}\|^{2}_{\mathcal{X}}$. Let $O_{0}=\mathbf{0}$ and let $O_{\ell}$ denote the operator obtained from the $\ell$-th operator-valued gradient descent sequence of $L$ with respect to $\|\cdot\|_{\mathcal{O}}$ as defined in Equation 22. Then there exist scalar step sizes $r_{0}^{\prime},\ldots,r_{m}^{\prime}$ such that if, for an $m$-layer continuum transformer $\mathcal{T}:=\mathcal{T}_{m}\circ...\circ\mathcal{T}_{0}$, $\widetilde{H}(U,W)]_{i,j}=\kappa(u^{(i)},w^{(j)})$, $\mathcal{W}_{v,\ell}=\begin{pmatrix}0&0\\ 0&-r^{\prime}_{\ell}I\end{pmatrix}$, $\mathcal{W}_{q,\ell}=I$, and $\mathcal{W}_{k,\ell}=I$ for each $\ell=0,...,m$, then for any $f\in\mathcal{X}$,
@@
From calculations in subsection Section D.1, we know that the $\ell$-th step of gradient descent has the form
@@
for all $j=0,\ldots,n$. In other words, $u^{(j)}-u^{(j)}_{\ell}$ is equal to the predicted label for $f$, if $f^{(j)}=f$”.
@@
Here, the first line follows from plugging in $\mathcal{W}_{q},\mathcal{W}_{v},\mathcal{W}_{k}$ in Appendix E. The second line follows from Equation 28 and the induction hypothesis. ∎
Appendix F Best Linear Unbiased Predictor
F.1 BLUP Coincides with Bayes Optimal
@@
F.2 In-Context Learning BLUP Proof
Proposition F.2.
Let $F=[f^{(1)},\ldots,f^{(n+1)}],U=[u^{(1)},\ldots,u^{(n+1)}]$. Let $\kappa:\mathcal{X}\times\mathcal{X}\to\mathcal{L}(\mathcal{X})$ be an operator-valued kernel. Assume that $U|F$ is a $\kappa$ Gaussian random variable per Definition 3.2. Let the activation function $\tilde{H}$ of the attention layer be defined as $\tilde{H}(U,W)]_{i,j}:=\kappa(u^{(i)},w^{(j)})$. Consider the operator gradient descent in Theorem 3.1. Then as the number of layers $m\to\infty$, the continuum transformer’s prediction at layer $m$ approaches the Best Linear Unbiased Predictor that minimizes the in-context loss in Equation 9.
@@ F.2 In-Context Learning BLUP Proof
From the premise of 3.1, we know that $\mathcal{W}_{q,\ell}=I,\mathcal{W}_{k,\ell}=I,\mathcal{W}_{v,\ell}=\begin{bmatrix}0&0\\ 0&-r^{\prime}_{\ell}I\end{bmatrix}$. Set $r_{\ell}=\delta I$ for all $\ell$, where $\delta$ is a positive constant satisfying $\|I-\delta\hat{\mathbb{K}}\|<1$, where the norm is the operator norm. From previous calculations done in Appendix E, we have:
@@
Again, using Appendix E, we have:
@@
Appendix G Optimization Convergence Proof
@@
Theorem G.5.
Suppose G.3 and G.4 hold. Let $f(r,\mathcal{W}_{q},\mathcal{W}_{k}):=L\left(\mathcal{W}_{v,\ell}=\left\{\begin{bmatrix}0&0\\ 0&r_{\ell}\end{bmatrix}\right\}_{\ell=0,\ldots,m},\mathcal{W}_{q,\ell},\mathcal{W}_{k,\ell}\right)$, where $L$ is as defined in Equation 9. Let $\mathcal{S}\subset\mathcal{O}^{m+1}\times\mathcal{O}^{m+1}$ denote the set of $(\mathcal{W}_{q},\mathcal{W}_{k})$ operators with the property that $(\mathcal{W}_{q},\mathcal{W}_{v})\in\mathcal{S}$ if and only if for all $\ell\in\{0,\ldots,m\}$, there exist scalars $b_{\ell},c_{\ell}\in\mathbb{R}$ such that $\mathcal{W}_{q,\ell}=b_{\ell}\Sigma^{-1/2}$ and $\mathcal{W}_{k,\ell}=c_{\ell}\Sigma^{-1/2}$. Then
@@
Clearly, Equation 31 $\sum_{\ell=0}^{k}(\partial_{r_{\ell}}L(r(t),\mathcal{W}_{q}(t),\mathcal{W}_{k}(t)))^{2}$. From Proposition G.6, we note that
@@
Proposition G.6.
Suppose $F,U$ satisfy G.3 and $\tilde{H}$ satisfies G.4. Suppose $\mathcal{W}_{q},\mathcal{W}_{k}\in\mathcal{S}$. Fix a layer $j\in\{0,\ldots,m\}$. For any $R\in\mathcal{L}(\mathcal{X})$, let $\mathcal{W}_{q}(R,j)$ denote the collection of operators where $\mathcal{W}_{q}(R,j)]_{j}:=\mathcal{W}_{q,j}+R$ and $\mathcal{W}_{q}(R,j)]_{\ell}=\mathcal{W}_{q,\ell}$ for all $\ell\neq j$. Take an arbitrary $R\in\mathcal{L}(\mathcal{X})$. Let
@@
where $R\Sigma^{1/2}$ denotes composition of operators and $Tr$ is the trace of an operator as defined in Definition G.10. Then for any $j\in\{0,\ldots,m\}$,
@@
Suppose $\mathcal{W}_{q}(R)$ be the collection of operators where $\mathcal{W}_{q}(R)]_{\ell}=\mathcal{W}_{q,\ell}+R$ for all $\ell\in\{0,\ldots,m\}$. Since Proposition G.6 holds for any $j$, and
@@
an immediate consequence of Proposition G.6 is that
@@
Suppose $\bar{Z}_{0}$ is equal to $Z_{0}$ everywhere, except the $(2,n+1)^{th}$ entry, where it is $c$ instead of 0. We compare the dynamics of $Z_{\ell}$ and $\bar{Z}_{\ell}$. Since $F_{0}=\bar{F}_{0}$, we see that $F_{\ell}=\bar{F}_{\ell}$. We claim that $\bar{U}_{\ell}-U_{\ell}=[0,\ldots,0,c]$. We prove this by induction. This is trivially true for $\ell=0$. Suppose this holds for step $\ell$, then at step $\ell+1$:
@@
Replacing $c$ by $u^{(n+1)}$ tells us that $\bar{Z}_{m+1}]_{2,n+1}=[Z_{m+1}]_{2,n+1}+u^{(n+1)}$. Hence, the loss in Equation 9 can be re-written as
@@
Since $A_{\ell}=0$, an immediate corollary of Appendix G is as follows.
Corollary G.8.
@@
The rest of the proof for Proposition G.6 has the following outline:
@@
Using Lemma G.11, we can write the in-context loss in Appendix G as
@@
Since (IM)(I-M) is a self-adjoint operator,
@@
where the last equality follows from G.3 and the linearity and cyclic property of trace.
Let $\mathcal{U}$ be a uniformly randomly sampled unitary operator. Let $\mathcal{U}_{\Sigma}=\Sigma^{1/2}\mathcal{U}\Sigma^{-1/2}$. Using G.3, $\Sigma^{1/2}\mathcal{M}\Sigma^{-1/2}F\stackrel{{\scriptstyle d}}{{=}}F$, we see that
@@
where the last equality uses the assumption $\mathbb{K}(\mathcal{U}_{\Sigma}F_{0})=\mathbb{K}(F_{0})$ from G.3.
@@
where the last equality follows from the rotational invariance of $\tilde{H}$ from G.4.
@@
where the first equality follows from Equation 41 and Equation 42 and the second equality follows from G.4.
@@
where the second equality follows from Equation 41, G.4 and Appendix G.
@@
Continuing from Appendix G, we get
@@
Appendix H Experiment Kernel Definitions
We provide below the explicit definitions of the kernels studied in the experiments of Section 5.1:
Kernels for $k_{x}(f,f^{\prime})$
Name Definition
Linear $\langle f,f^{\prime}\rangle_{\mathcal{X}}$
Kernels for $k_{y}(x,y)$
Name Definition
Laplace $\exp\left(-\frac{\|x-y\|_{2}}{\sigma_{y}}\right)$
Gaussian $\exp\left(-\frac{\|x-y\|_{2}^{2}}{2\sigma_{y}^{2}}\right)$
Appendix I BLUP Additional Experimental Results
@@
Appendix J Optimization Experiment Details
J.1 Optimization Kernel Assumptions
@@
J.2 Optimization Experiment Hyperparameters
@@
Appendix K Notation Reference
We provide below a comprehensive presentation of the notation from the paper.
Table 2: Notation for the data
Notation Description
$\mathcal{X}$ Hilbert space in which the input and output functions lie.
$f^{(i)},u^{(i)}$ $i$-th pair of input and output functions.
$Z_{\ell}$ A matrix in $\mathcal{X}^{2\times(n+1)}$ whose first row consists of input functions $f^{(i)};i=1,\ldots,n+1$, and whose second row consists of output functions $u^{(1)},\ldots,u^{(n)},0$. The zero is a placeholder for the predicted output of $f^{(n+1)}$ which will change as $Z_{\ell}$ passes through the transformer.
$X_{\ell}$ The first row of $Z_{\ell}$.
$\mathcal{O}$ Space of operators in which the true operator lies. This is assumed to be an RKHS.
$\mathbb{K}(F)$ Covariance operator of the output functions $U$ conditioned on the input functions $F$.
Table 3: Notation for the transformer architecture
Notation Description
$\mathcal{T}_{\ell}$ The $\ell$-th layer of the transformer.
$\mathcal{W}_{k,\ell},\mathcal{W}_{q,\ell},\mathcal{W}_{v,\ell}$ Key, query and value operators at the $\ell$-th layer respectively.
$\mathcal{M}$ Mask operator.
[…]
$r_{\ell}$ Scalars parametrizing the value operator at layer $\ell$.
Table 4: Notation for ICL, gradient descent, and the RKHS framework
Notation Description
$L(\cdot)$ In-context loss.
$O_{\ell}$ Operator at the $\ell$-th gradient descent step.
$\kappa$ Operator-valued kernel which takes an element in $\mathcal{X}\times\mathcal{X}$ to a bounded linear operator on $\mathcal{X}$.